# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 5.1 A lattice theoretical lemma
First we show $(V,\leq)$ is bowtie free. We verify the assumptions of Lemma 3.24. Assumption 1 is a direct consequence of Assumptions 2 and 3 of Theorem 3.27 (note that if $x$ is of type $\hat{s}_{1}$, then $V_{\geq}x$ is exactly the vertex set of $\operatorname{lk}(x,X)$). Let $x_{1}y_{1}x_{2}y_{2}$ be a 4-cycle in $X$ as in Assumption 2 of Lemma 3.24. Let $Y$ be the Helly graph as defined in Theorem 3.27. As any Helly graph satisfies the 4-wheel condition (see e.g. [CCHO14, Proposition 3.25]), there exists $z\in Y$ such that $z$ is adjacent to each of $\{x_{1},y_{1},x_{2},y_{2}\}$ in $Y$. As $x_{1},x_{2}$ are of type $\hat{s}_{1}$ and $y_{1},y_{2}$ are of type $\hat{s}_{n}$, by the definition of edges in $Y$ we know $z$ is also adjacent to each of $\{x_{1},y_{1},x_{2},y_{2}\}$ in $X$, as required.
@@
Suppose $\Lambda$ is a Dynkin diagram of type $\widetilde{C}_{n}$ with consecutive vertices being $\{s_{i}\}_{i=1}^{n}$. Then the vertex set of $\Delta_{\Lambda}$, endowed with the order from $s_{1}<s_{2}<\cdots<s_{n}$, is a bowtie free and flag poset.
@@ 5.2 An injective simplicial complex for type $A_{n}$
In the rest of this subsection, $\Lambda$ will be the Dynkin diagram of type $A_{n}$. Let consecutive vertices in $\Lambda$ be $\{s_{1},\ldots,s_{n}\}$. Let $\mathcal{A}$ be the reflection arrangement of type $A_{n}$. Up to a linear transformation, elements in $\mathcal{A}$ are $x_{i}=0$ for $1\leq i\leq n$, and $x_{i}=x_{j}$ for $1\leq i\neq j\leq n$. Let $H$ be the hyperplane $x_{1}=0$. Then a simple calculation implies that deconing of $\mathcal{A}$ with respect to $H$ gives the affine arrangement $\mathcal{B}$ in $\mathbb{R}^{n-1}$ made of $y_{i}=1$ for $11\leq i\leq n-$ 1, $y_{i}=0$ for $11\leq i\leq n-$ 1 and $y_{i}=y_{j}$ for $11\leq i\neq j\leq n-$ 1.
@@
Note that $\Sigma_{\mathcal{B}}$ can be identified as a maximal subcomplex of $\Sigma_{\mathcal{A}}$ contained in one side of $H$. This embedding $\Sigma_{\mathcal{B}}\to\Sigma_{\mathcal{A}}$ is dual to $D_{\mathcal{B}}\to S_{\mathcal{A}}$. We record the following consequence of Lemma 2.13.
@@
Let $\mathbb{D}_{\mathcal{B}}$ and $\mathbb{S}\mathbb{D}_{\mathcal{A}}$ be the Falk complex and spherical Deligne complex defined in Section 2.4. A vertex of $\mathbb{D}_{\mathcal{B}}$ is of type $\hat{s}_{i}$ if it maps to a vertex of type $\hat{s}_{i}$ under $\mathbb{D}_{\mathcal{B}}\to D_{\mathcal{B}}$. Note that $\mathbb{D}_{\mathcal{B}}$ is a simplicial complex of type $S=\{s_{1},\ldots,s_{n}\}$.
@@
The vertex set of $\mathbb{D}_{\mathcal{B}}$, endowed with the relation $<$ induced from $s_{1}<s_{2}<\cdots<s_{n}$, is a poset satisfying all the assumptions of Theorem 3.27.
@@
Note that $\mathbb{D}_{\mathcal{B}}$ is isomorphic to the inverse image of $D_{\mathcal{B}}$ (viewed as a subcomplex of $S_{\mathcal{A}}$) under the map $\mathbb{S}\mathbb{D}_{\mathcal{A}}\to S_{\mathcal{A}}$. Since the vertex set of $\mathbb{S}\mathbb{D}_{\mathcal{A}}$ (which is the Artin complex of $A_{n}$) with the induced order is a poset, so is the vertex set of $\mathbb{D}_{\mathcal{B}}$.
As each vertex of $\mathbb{D}_{\mathcal{B}}$ is lower bounded by a vertex of type $\hat{s}_{1}$ and upper bounded by a vertex of type $\hat{s}_{n}$, it suffices to verify Theorem 3.27 (3) for type $\hat{s}_{1}$ vertices and Theorem 3.27 (4) for type $\hat{s}_{n}$ vertices. Take a vertex $x$ of $\mathbb{D}_{\mathcal{B}}$ of type $s_{1}$. Then $x$ maps to $\bar{x}=(0,0,\ldots,0)$ under $\mathbb{D}_{\mathcal{B}}\to D_{\mathcal{B}}$. Note that the local arrangement of $\mathcal{B}$ at $\bar{c}$ (as defined in Section 2.4) is an arrangement of type $A_{n-1}$. By Lemma 5.5 below and Lemma 2.12, $\operatorname{lk}(x,\mathbb{D}_{\mathcal{B}})$ is type isomorphic to an Artin complex of type $B_{n-1}$, and we are done by Theorem 3.22. The verification of Theorem 3.27 (4) for type $\hat{s}_{n}$ vertices is similar. ∎
Now we consider the Artin group of type $B_{n}$ with consecutive vertices in the Dynkin diagram being $\{s_{1},\ldots,s_{n}\}$ and the edge between $s_{n-1}$ and $s_{n}$ is labeled by $4$. Consider the reflection arrangement $\mathcal{A}^{\prime}$ of type $B_{n}$ made of the following hyperplanes: $x_{i}=0$ for $1\leq i\leq n$, and $x_{i}\pm x_{j}=0$ for $1\leq i\neq j\leq n$. Then vertices of the complexes $S_{\mathcal{A}^{\prime}}$ and $\mathbb{S}\mathbb{D}_{\mathcal{A}^{\prime}}$ (cf. Section 2.4) are labeled by $\hat{s}_{i}$ for $1\leq i\leq n$.
@@
Let $S^{\prime}_{\mathcal{A}}$ be the subcomplex of $S_{\mathcal{A}}$ made of points with non-negative coordinates. We assume the vertex of $S^{\prime}_{\mathcal{A}}$ in the first octant corresponding to the line $x_{1}=x_{2}=\cdots=x_{n}$ is of type $\hat{s}_{n}$. Let $\mathbb{S}\mathbb{D}^{\prime}_{\mathcal{A}}$ be the inverse image of $S^{\prime}_{\mathcal{A}}$ under the map $\mathbb{S}\mathbb{D}_{\mathcal{A}}\to S_{\mathcal{A}}$. Then there is a simplicial isomorphism between $\mathbb{S}\mathbb{D}^{\prime}_{\mathcal{A}}$ and $\mathbb{S}\mathbb{D}_{\mathcal{A}^{\prime}}$ which preserve types of vertices.
@@
For the proof, we need an alternative description of $\mathbb{S}\mathbb{D}_{\mathcal{A}}$ and $\mathbb{S}\mathbb{D}_{\mathcal{A}^{\prime}}$, due to Allcock [All13, Theorem 7.3]. Let $M(\mathcal{A})=\mathbb{C}^{n}\setminus(\cup_{H\in\mathcal{A}}H\otimes\mathbb{C})$ be the complement of complexified arrangement associated with $\mathcal{A}$. Similarly we define $M(\mathcal{A}^{\prime})$. We endow $M(\mathcal{A})$ with the Euclidean metric induced from $\mathbb{C}^{n}$, which makes $M(\mathcal{A})$ an incomplete metric space whose metric completion (denoted $\bar{M}(\mathcal{A})$) is $\mathbb{C}^{n}$. Let $\bar{M}_{\mathbb{R}}(\mathcal{A})$ be the real part of $\bar{M}(\mathcal{A})\cong\mathbb{C}^{n}$. Let $\widetilde{M}(\mathcal{A})$ be the universal cover of $M(\mathcal{A})$, with the induced length metric from $M(\mathcal{A})$; and let $\widehat{M}(\mathcal{A})$ be the metric completion of $\widetilde{M}(\mathcal{A})$. Note that the covering map $\widetilde{M}(\mathcal{A})\to M(\mathcal{A})$ induced a map $p_{\mathcal{A}}:\widehat{M}(\mathcal{A})\to\bar{M}(\mathcal{A})$. We view $S_{\mathcal{A}}$ as the unit sphere in $\bar{M}_{\mathbb{R}}(\mathcal{A})$, hence $S_{\mathcal{A}}$ is a subset of $\bar{M}(\mathcal{A})$. Then $\mathbb{S}\mathbb{D}_{\mathcal{A}}$ is isomorphic to $p^{-1}_{\mathcal{A}}(S_{\mathcal{A}})$. A similar discussion applies if we replace $\mathcal{A}$ by $\mathcal{A}^{\prime}$. There is a finite sheeted covering map $f:M(\mathcal{A}^{\prime})\to M(\mathcal{A})$ induced by
@@
see [All02, Section 4]. Thus $\widetilde{M}(\mathcal{A}^{\prime})=\widetilde{M}(\mathcal{A})$ and $\widehat{M}(\mathcal{A}^{\prime})=\widehat{M}(\mathcal{A})$. This gives a homeomorphism $\widetilde{f}:\widetilde{M}(\mathcal{A}^{\prime})\to\widetilde{M}(\mathcal{A})$ with the following diagram commutes:
@@
As there is $L>0$ such that each point in $M(\mathcal{A}^{\prime})$ has an open neighborhood where $f$ restricts to a $L$-biLipschitz map, thus $\widetilde{f}$ is a biLipschitz homeomorphism, hence it extends to a biLipschitz homeomorphism $\widehat{f}:\widehat{M}(\mathcal{A}^{\prime})\to\widehat{M}(\mathcal{A})$ which fits into the following commutative diagram:
@@
$\mathbb{S}\mathbb{D}_{\mathcal{A}^{\prime}}\cong p^{-1}_{\mathcal{A}^{\prime}}(S_{\mathcal{A}^{\prime}})\cong p^{-1}_{\mathcal{A}}(K)\cong p^{-1}_{\mathcal{A}}(S^{\prime}_{\mathcal{A}})\cong\mathbb{S}\mathbb{D}^{\prime}_{\mathcal{A}}.$
@@
Suppose $\Lambda$ is a Dynkin diagram of type $A_{n}$. Let $\Lambda^{\prime}\subset\Lambda$ be a linear subgraph made of three vertices. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ is weakly flag.
@@
In the following proof, we assume $V\Delta_{\Lambda}$ is endowed with the order induced from $s_{1}<s_{2}<\cdots<s_{n}$. Assume vertices of $\Lambda^{\prime}$ are $t_{1}=s_{i},t_{2}=s_{i+1}$ and $t_{3}=s_{i+3}$. We will only prove upward weakly flagness here, as the proof of downward weakly flagness is almost identical. Take vertices $\{x_{i}\}_{i=1}^{3}$ in $\Delta_{\Lambda,\Lambda^{\prime}}$ of type $\hat{t}_{1}$ and $\{y_{i}\}_{i=1}^{3}$ of type $\hat{t}_{2}$. Suppose $y_{i}$ is a common upper bound for $x_{i}$ and $x_{i+1}$. Then we need to show $\{x_{1},x_{2},x_{3}\}$ have a common upper bound in $\Delta_{\Lambda,\Lambda^{\prime}}$.
Let $\omega$ be the 6-cycle $x_{1}y_{1}x_{2}y_{2}x_{3}y_{3}$ in $\Delta_{\Lambda,\Lambda^{\prime}}$. Let $\pi:\Delta_{\Lambda}\to\mathsf{C}_{\Lambda}\cong S_{\mathcal{A}}$ be the map induced by quotienting $\Delta_{\Lambda}$ by the action of the pure Artin group. By [Hua24, Proposition 3.1], it suffices to consider the case when $\pi(\omega)$ is a single edge.
We can assume without loss of generality that $\pi(\omega)\subset D_{\mathcal{B}}\subset S_{\mathcal{A}}$. As $\mathbb{D}_{\mathcal{B}}$ can be identified with subcomplex of $\mathbb{S}\mathbb{D}_{\mathcal{A}}\cong\Delta_{\Lambda}$ which is the inverse image of $D_{\mathcal{B}}$ under the map $\mathbb{S}\mathbb{D}_{\mathcal{A}}\to S_{\mathcal{A}}$, we know $\omega\subset\mathbb{D}_{\mathcal{B}}$. By Proposition 5.4 and Lemma 5.1, there is a vertex $z\in\mathbb{D}_{\mathcal{B}}$ which is the common upper bound for $\{x_{1},x_{2},x_{3}\}$ in $(V\mathbb{D}_{\mathcal{B}},\leq)$. As $(V\mathbb{D}_{\mathcal{B}},\leq)$ is bowtie free, we know $y_{i}$ is the join of $x_{i}$ and $x_{i+1}$. Thus $y_{i}<z$ in $(V\mathbb{D}_{\mathcal{B}},\leq)$. Then $z$ is of type $s_{j}$ with $j\geq i+2$. If $j=i+2$, then $z\in\Delta_{\Lambda,\Lambda^{\prime}}$ and we are done. Now assume $j>i+2$. Let $\operatorname{lk}^{-}(z,\Delta_{\Lambda})$ be the full subcomplex of $\operatorname{lk}(z,\Delta_{\Lambda})$ spanned by vertices $<z$. Then $\omega\subset\operatorname{lk}^{-}(z,\Delta_{\Lambda})$. Note that $\operatorname{lk}^{-}(z,\Delta_{\Lambda})\cong\Delta_{\Lambda_{j}}$ where $\Lambda_{j}$ is a Dynkin diagram of type $A_{j-1}$. By the same argument as before, we find $z^{\prime}\in\operatorname{lk}^{-}(z,\Delta_{\Lambda})$ such that $\omega\subset\operatorname{lk}^{-}(z^{\prime},\Delta_{\Lambda})$. Note that $z^{\prime}<z$. Repeating this procedure finitely many times will eventually give $z_{0}\in\Delta_{\Lambda}$ of type $\hat{t}_{3}$ such that $\omega\subset\operatorname{lk}^{-}(z_{0},\Delta_{\Lambda})$. Then $z_{0}\in\Delta_{\Lambda,\Lambda^{\prime}}$, as desired. ∎
@@
Proposition 5.7.
Suppose $\Lambda$ is a Dynkin diagram of type $A_{n}$. Let $\Lambda^{\prime}\subset\Lambda$ be a linear subgraph made of three vertices. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ with the induced Moussong metric from $\Delta_{\Lambda}$ is CAT$(1)$.

### v3

@@ 5.1 A lattice theoretical lemma
First we show $(V,\leq)$ is bowtie free. We verify the assumptions of Lemma 8.1. Assumption 1 is a direct consequence of Assumptions 2 and 3 of Theorem 2.18 (note that if $x$ is of type $\hat{s}_{1}$, then $V_{\geq}x$ is exactly the vertex set of $\operatorname{lk}(x,X)$). Let $x_{1}y_{1}x_{2}y_{2}$ be a 4-cycle in $X$ as in Assumption 2 of Lemma 8.1. Let $Y$ be the Helly graph as defined in Theorem 2.18. As any Helly graph satisfies the 4-wheel condition (see e.g. [CCHO14, Proposition 3.25]), there exists $z\in Y$ such that $z$ is adjacent to each of $\{x_{1},y_{1},x_{2},y_{2}\}$ in $Y$. As $x_{1},x_{2}$ are of type $\hat{s}_{1}$ and $y_{1},y_{2}$ are of type $\hat{s}_{n}$, by the definition of edges in $Y$ we know $z$ is also adjacent to each of $\{x_{1},y_{1},x_{2},y_{2}\}$ in $X$.
@@
Suppose $\Lambda$ is a Coxeter diagram of type $\widetilde{C}_{n}$ with consecutive nodes being $\{s_{i}\}_{i=1}^{n}$. Then the vertex set of $\Delta_{\Lambda}$, endowed with the order from $s_{1}<s_{2}<\cdots<s_{n}$, is a bowtie free and flag poset.
@@ 5.2 An injective simplicial complex for type $A_{n}$
In the rest of this subsection, $\Lambda$ will be the Coxeter diagram of type $A_{n}$. Let consecutive nodes in $\Lambda$ be $\{s_{1},\ldots,s_{n}\}$. Let $\mathcal{A}$ be the reflection arrangement of type $A_{n}$. Up to a linear transformation, elements in $\mathcal{A}$ are $x_{i}=0$ for $1\leq i\leq n$, and $x_{i}=x_{j}$ for $1\leq i\neq j\leq n$. Let $H$ be the hyperplane $x_{1}=0$. Then a simple calculation implies that deconing of $\mathcal{A}$ with respect to $H$ gives the affine arrangement $\mathcal{B}$ in $\mathbb{R}^{n-1}$ made of $y_{i}=1$ for $11\leq i\leq n-$ 1, $y_{i}=0$ for $11\leq i\leq n-$ 1 and $y_{i}=y_{j}$ for $11\leq i\neq j\leq n-$ 1.
@@
Let $\Sigma_{\mathcal{B}}$ and $\Sigma_{\mathcal{A}}$ be as in Section 4.1. Note that $\Sigma_{\mathcal{B}}$ can be identified as a maximal subcomplex of $\Sigma_{\mathcal{A}}$ contained in one side of $H$. This embedding $\Sigma_{\mathcal{B}}\to\Sigma_{\mathcal{A}}$ is dual to $D_{\mathcal{B}}\to S_{\mathcal{A}}$. We record the following consequence of Lemma 4.11.
@@
Let $\mathbb{D}_{\mathcal{B}}$ and $\mathbb{S}_{\mathcal{A}}$ be the Falk complex and spherical Deligne complex defined in Section 4.3. As $\widehat{\Sigma}_{\mathcal{B}}\to\widehat{\Sigma}_{\mathcal{A}}$ is $\pi_{1}$-injective, by the description of $\mathbb{D}_{\mathcal{B}}$ and $\mathbb{S}_{\mathcal{A}}$ in terms of certain subcomplexes of the universal covers of $\widehat{\Sigma}_{\mathcal{B}}$ and $\widehat{\Sigma}_{\mathcal{A}}$ in Section 4.3, we know there is an embedding $\mathbb{D}_{\mathcal{A}}\to\mathbb{S}_{\mathcal{A}}$ which is equivariant with respect to $\pi_{1}\widehat{\Sigma}_{\mathcal{B}}\curvearrowright\mathbb{D}_{\mathcal{B}}$, $\pi_{1}\widehat{\Sigma}_{\mathcal{A}}\curvearrowright\mathbb{S}_{\mathcal{A}}$ and $\pi_{1}\widehat{\Sigma}_{\mathcal{B}}\hookrightarrow\pi_{1}\widehat{\Sigma}_{\mathcal{A}}$. Hence we can view $\mathbb{D}_{\mathcal{A}}$ as a subcomplex of $\mathbb{S}_{\mathcal{A}}$, moreover, this subcomplex can be alternatively described as a connected component of the inverse image of $D_{\mathcal{B}}$ (viewed as a subcomplex of $S_{\mathcal{A}}$) under the map $\mathbb{S}_{\mathcal{A}}\to S_{\mathcal{A}}$.
A vertex of $\mathbb{D}_{\mathcal{B}}$ is of type $\hat{s}_{i}$ if it maps to a vertex of type $\hat{s}_{i}$ under $\mathbb{D}_{\mathcal{B}}\to D_{\mathcal{B}}$. Similarly, we define types of vertices in $\mathbb{S}_{\mathcal{A}}$. Note that $\mathbb{D}_{\mathcal{B}}$ and $\mathbb{S}_{\mathcal{A}}$ are simplicial complexes of type $S=\{s_{1},\ldots,s_{n}\}$. Moreover, $\mathbb{S}_{\mathcal{A}}$ is isomorphic to the Artin complex of type $A_{n}$ which preserves the types of vertices.
@@
The vertex set of $\mathbb{D}_{\mathcal{B}}$, endowed with the relation $<$ induced from $s_{1}<s_{2}<\cdots<s_{n}$ as in Definition 2.15, is a poset satisfying all the assumptions of Theorem 2.18.
@@
Since the vertex set of $\mathbb{S}_{\mathcal{A}}$ (which is the Artin complex of type $A_{n}$) with the induced order is a poset by Lemma 2.17, so is the vertex set of $\mathbb{D}_{\mathcal{B}}$.
As each vertex of $\mathbb{D}_{\mathcal{B}}$ is lower bounded by a vertex of type $\hat{s}_{1}$ and upper bounded by a vertex of type $\hat{s}_{n}$, it suffices to verify Theorem 2.18 (3) for type $\hat{s}_{1}$ vertices and Theorem 2.18 (4) for type $\hat{s}_{n}$ vertices. Take a vertex $x$ of $\mathbb{D}_{\mathcal{B}}$ of type $\hat{s}_{1}$. Then $x$ maps to $\bar{x}=(0,0,\ldots,0)$ under $\mathbb{D}_{\mathcal{B}}\to D_{\mathcal{B}}$. Note that the local arrangement of $\mathcal{B}$ at $\bar{x}$ (as defined in Section 4.3) is an arrangement of type $A_{n-1}$. By Lemma 5.5 below and Lemma 4.10, there is a vertex type preserving isomorphism between $\operatorname{lk}(x,\mathbb{D}_{\mathcal{B}})$ and the Artin complex of type $B_{n-1}$, and we are done by Theorem 2.37. The verification of Theorem 2.18 (4) for type $\hat{s}_{n}$ vertices is similar. ∎
Now we consider the Artin group of type $B_{n}$ with consecutive nodes in the Coxeter diagram being $\{s_{1},\ldots,s_{n}\}$ and the edge between $s_{n-1}$ and $s_{n}$ is labeled by $4$. Consider the reflection arrangement $\mathcal{A}^{\prime}$ of type $B_{n}$ made of the following hyperplanes: $x_{i}=0$ for $1\leq i\leq n$, and $x_{i}\pm x_{j}=0$ for $1\leq i\neq j\leq n$. Then vertices of the complexes $S_{\mathcal{A}^{\prime}}$ and $\mathbb{S}_{\mathcal{A}^{\prime}}$ (cf. Section 4.3) are labeled by $\hat{s}_{i}$ for $1\leq i\leq n$.
@@
Let $S^{\prime}_{\mathcal{A}}$ be the subcomplex of $S_{\mathcal{A}}$ made of points with non-negative coordinates. We assume the vertex of $S^{\prime}_{\mathcal{A}}$ in the first octant corresponding to the line $x_{1}=x_{2}=\cdots=x_{n}$ is of type $\hat{s}_{n}$. Let $\mathbb{S}^{\prime}_{\mathcal{A}}$ be the inverse image of $S^{\prime}_{\mathcal{A}}$ under the map $\mathbb{S}_{\mathcal{A}}\to S_{\mathcal{A}}$. Then there is a simplicial isomorphism between $\mathbb{S}^{\prime}_{\mathcal{A}}$ and $\mathbb{S}_{\mathcal{A}^{\prime}}$ which preserve types of vertices.
@@
For the proof, we need an alternative description of $\mathbb{S}_{\mathcal{A}}$ and $\mathbb{S}_{\mathcal{A}^{\prime}}$, due to Allcock [All13, Theorem 7.3]. Let $M(\mathcal{A})=\mathbb{C}^{n}\setminus(\cup_{H\in\mathcal{A}}H\otimes\mathbb{C})$ be the complement of complexified arrangement associated with $\mathcal{A}$. Similarly we define $M(\mathcal{A}^{\prime})$. We endow $M(\mathcal{A})$ with the Euclidean metric induced from $\mathbb{C}^{n}$, which makes $M(\mathcal{A})$ an incomplete metric space whose metric completion (denoted $\bar{M}(\mathcal{A})$) is $\mathbb{C}^{n}$. Let $\bar{M}_{\mathbb{R}}(\mathcal{A})$ be the real part of $\bar{M}(\mathcal{A})\cong\mathbb{C}^{n}$. Let $\widetilde{M}(\mathcal{A})$ be the universal cover of $M(\mathcal{A})$, with the induced length metric from $M(\mathcal{A})$; and let $\widehat{M}(\mathcal{A})$ be the metric completion of $\widetilde{M}(\mathcal{A})$. Note that the covering map $\widetilde{M}(\mathcal{A})\to M(\mathcal{A})$ induced a map $p_{\mathcal{A}}:\widehat{M}(\mathcal{A})\to\bar{M}(\mathcal{A})$. We view $S_{\mathcal{A}}$ as the unit sphere in $\bar{M}_{\mathbb{R}}(\mathcal{A})$, hence $S_{\mathcal{A}}$ is a subset of $\bar{M}(\mathcal{A})$. Then $\mathbb{S}_{\mathcal{A}}$ is isomorphic to $p^{-1}_{\mathcal{A}}(S_{\mathcal{A}})$. A similar discussion applies if we replace $\mathcal{A}$ by $\mathcal{A}^{\prime}$. There is a finite sheeted covering map $f:M(\mathcal{A}^{\prime})\to M(\mathcal{A})$ induced by
@@
see [All02, Section 4]. This gives a homeomorphism $\widetilde{f}:\widetilde{M}(\mathcal{A}^{\prime})\to\widetilde{M}(\mathcal{A})$ with the following diagram commutes:
@@
As there is $L>0$ such that each point in $M(\mathcal{A}^{\prime})$ has an open neighborhood where $f$ restricts to an $L$-biLipschitz map, thus the same property holds for $\widetilde{f}$. It follows that $\widetilde{f}$ is a biLipschitz homeomorphism with respect to the induced length metric, hence it extends to a biLipschitz homeomorphism $\widehat{f}:\widehat{M}(\mathcal{A}^{\prime})\to\widehat{M}(\mathcal{A})$ which fits into the following commutative diagram:
@@
$\mathbb{S}_{\mathcal{A}^{\prime}}\cong p^{-1}_{\mathcal{A}^{\prime}}(S_{\mathcal{A}^{\prime}})\cong p^{-1}_{\mathcal{A}}(K)\cong p^{-1}_{\mathcal{A}}(S^{\prime}_{\mathcal{A}})\cong\mathbb{S}^{\prime}_{\mathcal{A}}.$
@@
Suppose $\Lambda$ is a Coxeter diagram of type $A_{n}$. Let $\Lambda^{\prime}\subset\Lambda$ be a linear subgraph made of three nodes. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ is weakly flag.
@@
In the following proof, we assume $V\Delta_{\Lambda}$ is endowed with the order induced from $s_{1}<s_{2}<\cdots<s_{n}$. Assume nodes of $\Lambda^{\prime}$ are $t_{1}=s_{i},t_{2}=s_{i+1}$ and $t_{3}=s_{i+3}$. We will only prove upward weakly flagness here, as the proof of downward weakly flagness is almost identical. Take vertices $\{x_{i}\}_{i=1}^{3}$ in $\Delta_{\Lambda,\Lambda^{\prime}}$ of type $\hat{t}_{1}$ and $\{y_{i}\}_{i=1}^{3}$ of type $\hat{t}_{2}$. Suppose $y_{i}$ is a common upper bound for $x_{i}$ and $x_{i+1}$. Then we need to show $\{x_{1},x_{2},x_{3}\}$ have a common upper bound in $\Delta_{\Lambda,\Lambda^{\prime}}$.
Let $\omega$ be the 6-cycle $x_{1}y_{1}x_{2}y_{2}x_{3}y_{3}$ in $\Delta_{\Lambda,\Lambda^{\prime}}$. Let $\pi:\Delta_{\Lambda}\to\mathfrak{C}_{\Lambda}\cong S_{\mathcal{A}}$ be the map induced by quotienting $\Delta_{\Lambda}$ by the action of the pure Artin group. By [Hua24b, Proposition 3.1], it suffices to consider the case when $\pi(\omega)$ is a single edge.
We can assume without loss of generality that $\pi(\omega)\subset D_{\mathcal{B}}\subset S_{\mathcal{A}}$. As $\mathbb{D}_{\mathcal{B}}$ can be identified with the subcomplex of $\mathbb{S}_{\mathcal{A}}\cong\Delta_{\Lambda}$ which is a component of the inverse image of $D_{\mathcal{B}}$ under the map $\mathbb{S}_{\mathcal{A}}\to S_{\mathcal{A}}$, we can assume $\omega\subset\mathbb{D}_{\mathcal{B}}$. By Proposition 5.4 and Lemma 5.1, there is a vertex $z\in\mathbb{D}_{\mathcal{B}}$ which is the common upper bound for $\{x_{1},x_{2},x_{3}\}$ in $(V\mathbb{D}_{\mathcal{B}},\leq)$. As $(V\mathbb{D}_{\mathcal{B}},\leq)$ is bowtie free, we know $y_{i}$ is the join of $x_{i}$ and $x_{i+1}$. Thus $y_{i}<z$ in $(V\mathbb{D}_{\mathcal{B}},\leq)$. Then $z$ is of type $s_{j}$ with $j\geq i+2$. If $j=i+2$, then $z\in\Delta_{\Lambda,\Lambda^{\prime}}$ and we are done. Now assume $j>i+2$. Let $\operatorname{lk}^{-}(z,\Delta_{\Lambda})$ be the full subcomplex of $\operatorname{lk}(z,\Delta_{\Lambda})$ spanned by vertices $<z$. Then $\omega\subset\operatorname{lk}^{-}(z,\Delta_{\Lambda})$. Note that $\operatorname{lk}^{-}(z,\Delta_{\Lambda})\cong\Delta_{\Lambda_{j}}$ where $\Lambda_{j}$ is a Coxeter diagram of type $A_{j-1}$. By the same argument as before, we find $z^{\prime}\in\operatorname{lk}^{-}(z,\Delta_{\Lambda})$ such that $\omega\subset\operatorname{lk}^{-}(z^{\prime},\Delta_{\Lambda})$. Note that $z^{\prime}<z$. Repeating this procedure finitely many times will eventually give $z_{0}\in\Delta_{\Lambda}$ of type $\hat{t}_{3}$ such that $\omega\subset\operatorname{lk}^{-}(z_{0},\Delta_{\Lambda})$. Then $z_{0}\in\Delta_{\Lambda,\Lambda^{\prime}}$, as desired. ∎
@@
Corollary 5.7.
Suppose $\Lambda$ is a Coxeter diagram of type $A_{n}$. Let $\Lambda^{\prime}\subset\Lambda$ be a linear subgraph made of three nodes. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ with its simplices equipped with $A_{3}$-shape is CAT$(1)$.